# FoundTS: Comprehensive and Unified Benchmarking of Foundation Models for Time Series Forecasting

### v1

@@ Abstract
Time Series Forecasting (TSF) is core functionality in a multitude of applications, including in finance, weather services, and energy management Wu et al. (2024); Pan et al. (2023); Wan et al. (2022); Qin et al. (2023). Given historical observations, predicting future states is valuable for decision making and taking appropriate actions. Consequently, TSF is a very active research field, as evidenced by the continued emergence of prediction models. However, most existing TSF models require training on specific datasets in preparation for performing inference on corresponding datasets. Such models (called specific models in this paper to be distinguished from foundation models) do not generalize well and experience suboptimal performance when applied to new or unseen data Shi et al. (2024); Huang et al. (2023); Nie et al. (2022); Lin et al. (2024). Efforts to address challenges such as these have led to a recent surge in the development of foundation models for TSF Liang et al. (2024); Woo et al. (2024); Liu et al. (2024b); Zhou et al. (2023).
While foundation models encompassing diverse architectures and training paradigms continue to appear, our understanding of their strengths and limitations remains limited. Existing research on understanding such models has focused primarily on the qualitative analysis and categorization of foundation models for TSF Liang et al. (2024); Jin et al. (2023b; 2024). For instance,  Jin et al. (2024) propose a potential framework for LLM-based time series analysis, highlighting key opportunities and challenges for future research and advocating for increased interdisciplinary collaboration and exploration in this promising field. Similarly,  Liang et al. (2024) adopt a methodology-centered classification approach, outlining critical elements of time series foundation models, such as model architectures, pre-training techniques, and adaptation strategies. However, these studies often lack a quantitative evaluation of foundation models, which is crucial for assessing and comparing performance. Quantitative analysis not only provides a clearer understanding of model performance; it also enables researchers to make informed decisions about model selection and improvements.
@@
Further, proposals for foundation models often adopt different experimental setups, making it difficult to compare meaningfully the performance of different foundation models based only on the existing body of performance studies. As shown in Table 1, few-shot learning studies employ different types of sampling. Some use uniform window sampling Liu et al. (2024b); Gao et al. (2024), while others use front-end window sampling Jin et al. (2023a); Pan et al. (2024). Additionally, studies employ different lookback lengths and sampling ratios.
Robust and thorough benchmarks enable researchers to evaluate new models more rigorously, which is crucial for advancing the state-of-the-art Tan et al. (2020); Qiao et al. (2024). Most time series forecasting benchmarks target performance evaluation of specific models, while evaluations of foundation models are relatively scarce—see Table 2. The only benchmark targeting foundation models Zhang et al. (2023) has two notable limitations: 1) the types of foundation models considered is not sufficiently comprehensive, as LLM-based models are ignored; 2) the evaluation strategies supported are too limited to reflect fully the performance of foundation models, as few-shot settings are not considered.
@@ 2.1 Time series forecasting
Statistical learning models Box & Jenkins (1968); Hyndman et al. (2008), though theoretically robust, struggle to capture nonlinear trends, thus limiting their predictive accuracy. Machine learning models Chen & Guestrin (2016); Friedman (2001) are better at capturing non-linear relationships and complex patterns, but often require manual feature engineering and model design. Deep learning models Zhong et al. (2023); Dai et al. (2024); Wen et al. (2023); Wang et al. (2024a); Lin et al. (2023); Zhou et al. (2021) leverage the representation learning capabilities of deep neural networks on rich datasets, often outperforming the other two categories of techniques at predictive accuracy. However, all these models are limited by their strong coupling of training and inferencing data. They may not perform well on new or unseen data.
Foundation models for time series forecasting Liang et al. (2024) can be divided into two categories: LLM-based models and time series pre-trained models. Both exhibit outstanding zero-shot and few-shot prediction capabilities on unseen time series datasets. LLM-based models Zhou et al. (2023), with their vast language understanding and context processing cap abilities, are capable of high-quality forecasts when faced with unseen data. Time series pre-trained models Woo et al. (2024); Liu et al. (2024b), through pre-training on large time series datasets, exhibit generalization capabilities, enabling them to perform forecasting with limited training data.
@@
M3 Makridakis & Hibon (2000) $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
M4 Makridakis et al. (2018) $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
LTSF-Liner Zeng et al. (2023) $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
BasicTS Liang et al. (2022) $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
BasicTS+ Shao et al. (2023) $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
Monash Godahewa et al. (2021) $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
Libra Bauer et al. (2021) $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
ProbTS Zhang et al. (2023) $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\surd$
TSLib Wang et al. (2024d) $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
TFB Qiu et al. (2024) $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
@@ 2.2 Time Series Forecasting Benchmarks
Several benchmarks have recently been proposed for TSF—see Table 2. However, their inherent limitations make comprehensive and fair comparison of foundation models and specific models for time series forecasting out of reach. First, most benchmarks target only specific models and ignore time series foundation models. As shown in Table 2, the only exception is ProbTS Zhang et al. (2023), which, however, only cover time series pre-trained models, not LLM-based models. With foundation models now offering impressive features like zero-shot prediction, there is a need for fair and thorugh comparisons with specific models, considering the challenges they pose, such as high computational costs.
Second, current benchmarks do not support diverse evaluation strategies. Most time series forecasting benchmarks disregard emerging features like zero-shot and few-shot prediction, focusing instead on full-shot scenarios. While ProbTS Zhang et al. (2023) supports zero-shot prediction, it does not support few-shot prediction, which enables models to leverage small amounts of relevant data to fine-tune their performance. This capability not only enhances accuracy but also increases a model’s flexibility at adapting to new tasks, making it more effective in dynamic environments and across diverse environments. Furthermore, the absence of standardized sampling methods for few-shot prediction compromizes fair model comparison. There is a need for a more inclusive and fair evaluation pipeline.
@@ 3 FoundTS
(1) Domains: We include datasets from ten domains, including stock (NASDAQ Feng et al. (2019)), health (ILI Wu et al. (2021)), energy (Solar Lai et al. (2018)), electricity (ETT Zhou et al. (2021) and Electricity Trindade (2015)), environment (Weather Wu et al. (2021)), traffic (Traffic Wu et al. (2021)), nature (ZafNoo Poyatos et al. (2020)), banking (NN5 Taieb et al. (2012)), web (Wike2000 Gasthaus et al. (2019)), and economics (Exchange Lai et al. (2018)), for evaluation.
(2) Characteristics: We consider a range of important time series characteristics, including seasonality, trend, stationarity, transition, shifting, correlation, and non-Gaussianity Qiu et al. (2024); Zhang et al. (2023). Seasonality refers to repeating patterns or cycles at regular intervals. Trend indicates overall movements in a time series. Stationarity reflects the statistical properties of a time series, such as mean and variance, which do not change over time. Transition represents sudden or gradual shifts in a time series. Shifting refers to changes in the level or timing of the data and includes vertical and horizontal shifts. Correlation represents the relationship or dependence among different channels. Non-Gaussianity (N-Gau) represents deviations from normal distribution, often exhibiting skewness or kurtosis. The formula used to calculate these characteristics can be found in Appendix B.
@@ 3.2.1 Time Series Foundation Models
LLMs-based methods leverage the strong representational capacity and sequential modeling capability of LLMs to capture complex patterns for time series modeling. To more comprehensively evaluate the foundation models, we incorporate existing LLMs-based methods into FoundTS, focusing primarily on parameter-efficient fine-tuning and prompting. 1) Parameter-efficient fine-tuning methods: GPT4TS Zhou et al. (2023), S2IP-LLM Pan et al. (2024) selectively adjust specific parameters such as positional encoding and layer normalization of LLMs, enabling the model to quickly adapt to time series while retaining most of pre-trained knowledge. 2) Prompting methods, such as UniTime Liu et al. (2024a) and Time-LLM Jin et al. (2023a), focus on designing prompts, such as learnable prompts, prompt pools, and domain-specific instructions to activate time series knowledge in LLMs.
@@
Pre-training on multi-domain time series data has gained significant attention in recent years. We incorporate time series pre-trained models into FoundTS, categorizing them into four types based on the pre-training approach: reconstruction, autoregressive, direct prediction, and hybrid training. 1) Reconstruction methods: MOIRAI Woo et al. (2024), UniTS Gao et al. (2024), Moment Goswami et al. (2024) restore the features of time series data, enabling them to extract valuable information in an unsupervised manner. This type of method mainly adopts the encoder architecture. 2) Autoregressive methods: TimesFM Das et al. (2023), Timer Liu et al. (2024b), employ next token prediction to learn time series representation. This type of method mainly adopts the decoder architecture. 3) Direct prediction methods: TTM Ekambaram et al. (2024), unify the training process between pre-training and downstream tasks, allowing models to exhibit strong adaptability when transitioning to downstream forecasting tasks. 4) Hybrid pre-training methods: ROSE Wang et al. (2024c), combines the strengths of both reconstruction and direct prediction to learn generalized time series representations.
@@ 3.2.2 Time Series specific Models
Time series specific models typically require training on specific datasets and perform inference on the corresponding datasets. To better showcase the capabilities of time series foundation models, we select several SOTA specific models for comparison. We include: 1) CNN-based models: TimesNet Wu et al. (2022), which treat time series as sequences of vectors and leverage CNNs to capture temporal dependencies. 2) Transformer-based models: FEDformer Zhou et al. (2022), iTransformer Liu et al. (2023), and PatchTST Nie et al. (2022), which are capable of capturing more complex temporal dynamics, leading to significantly improved forecasting performance. 3) MLP-based models: FITS Xu et al. (2024), TimeMixer Wang et al. (2024b), and DLinear Zeng et al. (2023), with their simple architecture and relatively few parameters, have demonstrated strong forecasting accuracy as well.
@@ 3.3.2 Settings and metrics
(2) Standardized data splitting and loading: We standardize the division of the training, validation and test datasets, as well as the partitioning of each time series sample for all models. To ensure that different models use a consistent test length, we do not apply the “Drop Last” operation during testing Qiu et al. (2024).
@@ A.1 Datasets
We use the following 14 multivariate time-series datasets which cover 10 domains for forecasting: ETT Zhou et al. (2021) datasets contain 7 variates collected from two different electric transformers from July 2016 to July 2018. It consists of four subsets, of which ETTh1/ETTh2 are recorded hourly, and ETTm1/ETTm2 are recorded every 15 minutes. Electricity Trindade (2015) contains the electricity consumption of 321 customers from July 2016 to July 2019, recorded hourly. Traffic Wu et al. (2021) contains road occupancy rates measured by 862 sensors on freeways in the San Francisco Bay Area from 2015 to 2016, recorded hourly. Solar Lai et al. (2018) records solar power generation from 137 PV plants in 2006, every 10 minutes. Weather Wu et al. (2021) collects 21 meteorological indicators, such as temperature and barometric pressure, for Germany in 2020, recorded every 10 minutes. Exchange Lai et al. (2018) collects the daily exchange rates of eight countries. ZafNoo Poyatos et al. (2020) is collected from the Sapflux data project and includes sap flow measurements and environmental variables. ILI Wu et al. (2021) records indicators of patients data from Centers for Disease Control and Prevention. NASDAQ Feng et al. (2019) records opening price, closing price, trading volume, lowest price, and highest price. NN5 Taieb et al. (2012) is from banking and records the daily cash withdrawals from ATMs in the UK. Wike2000 Gasthaus et al. (2019) is the daily page views of 2000 Wikipedia pages. Table 8 lists statistics of the 14 multivariate time series datasets. Please note that the values for the seven characteristics—seasonality, trend, stationarity, transition, shifting, correlation, and non-Gaussianity—in the table are the results of min-max normalization.
@@ A.2 Time Series Forecasting Models
In the realm of time series forecasting, numerous models have surfaced in recent years. We choose models with superior predictive performance in our benchmark, including the pre-trained time series models: ROSE Wang et al. (2024c), TimesFM Das et al. (2023), Timer Liu et al. (2024b), TTM Ekambaram et al. (2024), Moirai Woo et al. (2024), and UniTS Gao et al. (2024); The LLM-based models: GPT4TS Zhou et al. (2023), S2IP-LLM Pan et al. (2024), UniTime Liu et al. (2024a) and Time-LLM Jin et al. (2023a); And the specific models: TimesNet Wu et al. (2022), Fedformer Zhou et al. (2022), iTransformer Liu et al. (2023), PatchTST Nie et al. (2022), FITS Xu et al. (2024), TimeMixer Wang et al. (2024b), and Dlinear Zeng et al. (2023). The specific descriptions for each of these models—see Table 9.

### v2

@@ Abstract
Time Series Forecasting (TSF) is core functionality in a multitude of applications, including in finance, weather services, and energy management (Wu et al., 2024; Pan et al., 2023; Wan et al., 2022; Qin et al., 2023). Given historical observations, predicting future states is valuable for decision making and taking appropriate actions. Consequently, TSF is a very active research field, as evidenced by the continued emergence of prediction models. However, most existing TSF models require training on specific datasets in preparation for performing inference on corresponding datasets. Such models (called specific models in this paper to be distinguished from foundation models) do not generalize well and experience suboptimal performance when applied to new or unseen data (Shi et al., 2024; Huang et al., 2023; Nie et al., 2022; Lin et al., 2024). Efforts to address challenges such as these have led to a recent surge in the development of foundation models for TSF (Liang et al., 2024; Woo et al., 2024; Liu et al., 2024b; Zhou et al., 2023).
While foundation models encompassing diverse architectures and training paradigms continue to appear, our understanding of their strengths and limitations remains limited. Existing research on understanding such models has focused primarily on the qualitative analysis and categorization of foundation models for TSF (Liang et al., 2024; Jin et al., 2023b; 2024). For instance,  (Jin et al., 2024) propose a potential framework for LLM-based time series analysis, highlighting key opportunities and challenges for future research and advocating for increased interdisciplinary collaboration and exploration in this promising field. Similarly,  (Liang et al., 2024) adopt a methodology-centered classification approach, outlining critical elements of time series foundation models, such as model architectures, pre-training techniques, and adaptation strategies. However, these studies often lack a quantitative evaluation of foundation models, which is crucial for assessing and comparing performance. Quantitative analysis not only provides a clearer understanding of model performance; it also enables researchers to make informed decisions about model selection and improvements.
@@
Further, proposals for foundation models often adopt different experimental setups, making it difficult to compare meaningfully the performance of different foundation models based only on the existing body of performance studies. As shown in Table 1, few-shot learning studies employ different types of sampling. Some use uniform window sampling (Liu et al., 2024b; Gao et al., 2024), while others use front-end window sampling (Jin et al., 2023a; Pan et al., 2024). Additionally, studies employ different lookback lengths and sampling ratios.
Robust and thorough benchmarks enable researchers to evaluate new models more rigorously, which is crucial for advancing the state-of-the-art (Tan et al., 2020; Qiao et al., 2024). Most time series forecasting benchmarks target performance evaluation of specific models, while evaluations of foundation models are relatively scarce—see Table 2. The only benchmark targeting foundation models (Zhang et al., 2023) has two notable limitations: 1) the types of foundation models considered is not sufficiently comprehensive, as LLM-based models are ignored; 2) the evaluation strategies supported are too limited to reflect fully the performance of foundation models, as few-shot settings are not considered.
@@ 2.1 Time series forecasting
Statistical learning models (Box & Jenkins, 1968; Hyndman et al., 2008), though theoretically robust, struggle to capture nonlinear trends, thus limiting their predictive accuracy. Machine learning models (Chen & Guestrin, 2016; Friedman, 2001) are better at capturing non-linear relationships and complex patterns, but often require manual feature engineering and model design. Deep learning models (Zhong et al., 2023; Dai et al., 2024; Wen et al., 2023; Wang et al., 2024a; Lin et al., 2023; Zhou et al., 2021) leverage the representation learning capabilities of deep neural networks on rich datasets, often outperforming the other two categories of techniques at predictive accuracy. However, all these models are limited by their strong coupling of training and inferencing data. They may not perform well on new or unseen data.
Foundation models for time series forecasting (Liang et al., 2024) can be divided into two categories: LLM-based models and time series pre-trained models. Both exhibit outstanding zero-shot and few-shot prediction capabilities on unseen time series datasets. LLM-based models (Zhou et al., 2023), with their vast language understanding and context processing cap abilities, are capable of high-quality forecasts when faced with unseen data. Time series pre-trained models (Woo et al., 2024; Liu et al., 2024b), through pre-training on large time series datasets, exhibit generalization capabilities, enabling them to perform forecasting with limited training data.
@@
M3 (Makridakis & Hibon, 2000) $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
M4 (Makridakis et al., 2018) $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
LTSF-Liner (Zeng et al., 2023) $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
BasicTS (Liang et al., 2022) $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
BasicTS+ (Shao et al., 2023) $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
Monash (Godahewa et al., 2021) $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
Libra (Bauer et al., 2021) $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
ProbTS (Zhang et al., 2023) $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\surd$
TSLib (Wang et al., 2024d) $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
TFB (Qiu et al., 2024) $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
@@ 2.2 Time Series Forecasting Benchmarks
Several benchmarks have recently been proposed for TSF—see Table 2. However, their inherent limitations make comprehensive and fair comparison of foundation models and specific models for time series forecasting out of reach. First, most benchmarks target only specific models and ignore time series foundation models. As shown in Table 2, the only exception is ProbTS (Zhang et al., 2023), which, however, only cover time series pre-trained models, not LLM-based models. With foundation models now offering impressive features like zero-shot prediction, there is a need for fair and thorugh comparisons with specific models, considering the challenges they pose, such as high computational costs.
Second, current benchmarks do not support diverse evaluation strategies. Most time series forecasting benchmarks disregard emerging features like zero-shot and few-shot prediction, focusing instead on full-shot scenarios. While ProbTS (Zhang et al., 2023) supports zero-shot prediction, it does not support few-shot prediction, which enables models to leverage small amounts of relevant data to fine-tune their performance. This capability not only enhances accuracy but also increases a model’s flexibility at adapting to new tasks, making it more effective in dynamic environments and across diverse environments. Furthermore, the absence of standardized sampling methods for few-shot prediction compromizes fair model comparison. There is a need for a more inclusive and fair evaluation pipeline.
@@ 3 FoundTS
(1) Domains: We include datasets from ten domains, including stock (NASDAQ (Feng et al., 2019)), health (ILI (Wu et al., 2021)), energy (Solar (Lai et al., 2018)), electricity (ETT (Zhou et al., 2021) and Electricity (Trindade, 2015)), environment (Weather (Wu et al., 2021)), traffic (Traffic (Wu et al., 2021)), nature (ZafNoo (Poyatos et al., 2020)), banking (NN5 (Taieb et al., 2012)), web (Wike2000 (Gasthaus et al., 2019)), and economics (Exchange (Lai et al., 2018)), for evaluation.
(2) Characteristics: We consider a range of important time series characteristics, including seasonality, trend, stationarity, transition, shifting, correlation, and non-Gaussianity (Qiu et al., 2024; Zhang et al., 2023). Seasonality refers to repeating patterns or cycles at regular intervals. Trend indicates overall movements in a time series. Stationarity reflects the statistical properties of a time series, such as mean and variance, which do not change over time. Transition represents sudden or gradual shifts in a time series. Shifting refers to changes in the level or timing of the data and includes vertical and horizontal shifts. Correlation represents the relationship or dependence among different channels. Non-Gaussianity (N-Gau) represents deviations from normal distribution, often exhibiting skewness or kurtosis. The formula used to calculate these characteristics can be found in Appendix B.
@@ 3.2.1 Time Series Foundation Models
LLMs-based methods leverage the strong representational capacity and sequential modeling capability of LLMs to capture complex patterns for time series modeling. To more comprehensively evaluate the foundation models, we incorporate existing LLMs-based methods into FoundTS, focusing primarily on parameter-efficient fine-tuning and prompting. 1) Parameter-efficient fine-tuning methods: GPT4TS (Zhou et al., 2023), S2IP-LLM (Pan et al., 2024) selectively adjust specific parameters such as positional encoding and layer normalization of LLMs, enabling the model to quickly adapt to time series while retaining most of pre-trained knowledge. 2) Prompting methods, such as UniTime (Liu et al., 2024a) and Time-LLM (Jin et al., 2023a), focus on designing prompts, such as learnable prompts, prompt pools, and domain-specific instructions to activate time series knowledge in LLMs.
@@
Pre-training on multi-domain time series data has gained significant attention in recent years. We incorporate time series pre-trained models into FoundTS, categorizing them into four types based on the pre-training approach: reconstruction, autoregressive, direct prediction, and hybrid training. 1) Reconstruction methods: MOIRAI (Woo et al., 2024), UniTS (Gao et al., 2024), Moment (Goswami et al., 2024) restore the features of time series data, enabling them to extract valuable information in an unsupervised manner. This type of method mainly adopts the encoder architecture. 2) Autoregressive methods: TimesFM (Das et al., 2023), Timer (Liu et al., 2024b), employ next token prediction to learn time series representation. This type of method mainly adopts the decoder architecture. 3) Direct prediction methods: TTM (Ekambaram et al., 2024), unify the training process between pre-training and downstream tasks, allowing models to exhibit strong adaptability when transitioning to downstream forecasting tasks. 4) Hybrid pre-training methods: ROSE (Wang et al., 2024c), combines the strengths of both reconstruction and direct prediction to learn generalized time series representations.
@@ 3.2.2 Time Series specific Models
Time series specific models typically require training on specific datasets and perform inference on the corresponding datasets. To better showcase the capabilities of time series foundation models, we select several SOTA specific models for comparison. We include: 1) CNN-based models: TimesNet (Wu et al., 2022), which treat time series as sequences of vectors and leverage CNNs to capture temporal dependencies. 2) Transformer-based models: FEDformer (Zhou et al., 2022), iTransformer (Liu et al., 2023), and PatchTST (Nie et al., 2022), which are capable of capturing more complex temporal dynamics, leading to significantly improved forecasting performance. 3) MLP-based models: FITS (Xu et al., 2024), TimeMixer (Wang et al., 2024b), and DLinear (Zeng et al., 2023), with their simple architecture and relatively few parameters, have demonstrated strong forecasting accuracy as well.
@@ 3.3.2 Settings and metrics
(2) Standardized data splitting and loading: We standardize the division of the training, validation and test datasets, as well as the partitioning of each time series sample for all models. To ensure that different models use a consistent test length, we do not apply the “Drop Last” operation during testing (Qiu et al., 2024).
@@ A.1 Datasets
We use the following 14 multivariate time-series datasets which cover 10 domains for forecasting: ETT (Zhou et al., 2021) datasets contain 7 variates collected from two different electric transformers from July 2016 to July 2018. It consists of four subsets, of which ETTh1/ETTh2 are recorded hourly, and ETTm1/ETTm2 are recorded every 15 minutes. Electricity (Trindade, 2015) contains the electricity consumption of 321 customers from July 2016 to July 2019, recorded hourly. Traffic (Wu et al., 2021) contains road occupancy rates measured by 862 sensors on freeways in the San Francisco Bay Area from 2015 to 2016, recorded hourly. Solar (Lai et al., 2018) records solar power generation from 137 PV plants in 2006, every 10 minutes. Weather (Wu et al., 2021) collects 21 meteorological indicators, such as temperature and barometric pressure, for Germany in 2020, recorded every 10 minutes. Exchange (Lai et al., 2018) collects the daily exchange rates of eight countries. ZafNoo (Poyatos et al., 2020) is collected from the Sapflux data project and includes sap flow measurements and environmental variables. ILI (Wu et al., 2021) records indicators of patients data from Centers for Disease Control and Prevention. NASDAQ (Feng et al., 2019) records opening price, closing price, trading volume, lowest price, and highest price. NN5 (Taieb et al., 2012) is from banking and records the daily cash withdrawals from ATMs in the UK. Wike2000 (Gasthaus et al., 2019) is the daily page views of 2000 Wikipedia pages. Table 8 lists statistics of the 14 multivariate time series datasets. Please note that the values for the seven characteristics—seasonality, trend, stationarity, transition, shifting, correlation, and non-Gaussianity—in the table are the results of min-max normalization.
@@ A.2 Time Series Forecasting Models
In the realm of time series forecasting, numerous models have surfaced in recent years. We choose models with superior predictive performance in our benchmark, including the pre-trained time series models: ROSE (Wang et al., 2024c), TimesFM (Das et al., 2023), Timer (Liu et al., 2024b), TTM (Ekambaram et al., 2024), Moirai (Woo et al., 2024), and UniTS (Gao et al., 2024); The LLM-based models: GPT4TS (Zhou et al., 2023), S2IP-LLM (Pan et al., 2024), UniTime (Liu et al., 2024a) and Time-LLM (Jin et al., 2023a); And the specific models: TimesNet (Wu et al., 2022), Fedformer (Zhou et al., 2022), iTransformer (Liu et al., 2023), PatchTST (Nie et al., 2022), FITS (Xu et al., 2024), TimeMixer (Wang et al., 2024b), and Dlinear (Zeng et al., 2023). The specific descriptions for each of these models—see Table 9.